# Training-free Long Video Generation with Chain of Diffusion Model Experts

### v2

@@ Introduction
Figure 1: Generated videos of our approach ConFiner-Long with only 9 inference steps.
@@
Therefore, based on ConFiner, we propose ConFiner-Long framework with three strategies to ensure the coherence and consistency between segments. As the initial noise significantly impacts the final videos, we first introduce a segments consistency initialization strategy to ensure the consistency of the initial noise between segments by sharing a base noise. Additionally, in order to enhance the consistency of the denoising direction between segments, we propose a consistency guidance strategy that uses the gradient of noise differences between two segments to guide the denoising direction. Also, to address the problem of flickering at the junctions of segments, we design a staggered refinement strategy that staggers the control stage and the refinement stage. It places the tail of one video structure and the head of the next into the same refinement process to achieve more natural transitions between segments.
Figure 2: Pipeline of Our ConFiner and ConFiner-Long. ConFiner decouples the video generation process. Firstly, control expert generates a video structure. Subsequently, temporal and spatial experts perform the refinement of spatio-temporal details. Spatial and temporal experts work together in refinement stage with our coordinated denoising. By adding consistency initialization, consistency guidance and staggered refinement to ConFiner, ConFiner-Long can generate coherent long videos.
@@
First, we design consistency initialization strategy to promote consistency between segments. The initial noise affects the structure of the generated video significantly. To improve the consistency between segments, we first sample a $Noise\_{base}\in\mathbf{R}^{H\times W\times C\times F}$, which is then subjected to frame-wise shuffling to obtain the initial noise for each segment. Sharing base noise enhances the consistency between segments while shuffling maintains a little randomness.
@@
Moreover, we propose a consistency guidance mechanism that promotes the generated content to follow the preceding segment. Specifically, during the sampling process of video segments, we use the gradient of the L2 loss to guide the sampling direction. The L2 loss is calculated between the predicted noise of the current segment and the noise in the previous segment. The guided noise is calculated as follows:
@@
Figure 3: Ablation Studys on Three Strategies of ConFiner-Long.
@@
Ablation Studys on Strategies of ConFiner-Long
In this section, we conducted ablation experiments on three strategies within the ConFiner-Long framework. Using the same preceding video segments, we generated subsequent video segments with either all strategies or only two. The visual comparison of the four video segments against the preceding one is shown in fig. 3.

### v3

@@ Introduction
Figure 1: Visual Comparison between Our ConFiner-Long and StreamingT2V.
@@
Therefore, based on ConFiner, we propose ConFiner-Long framework with three strategies to ensure the coherence and consistency between segments. As the initial noise significantly impacts the final videos, we first introduce a segments consistency initialization strategy to ensure the consistency of the initial noise between segments by sharing a base noise. Additionally, in order to enhance the coherence of the motion trend between segments, we propose a coherence guidance strategy that uses the gradient of noise differences between two segments to guide the denoising direction. Also, to address the problem of flickering at the junctions of segments, we design a staggered refinement strategy that staggers the control stage and the refinement stage. It places the tail of one video structure and the head of the next into the same refinement process to achieve more natural transitions between segments.
Figure 2: Pipeline of Our ConFiner and ConFiner-Long. ConFiner decouples the video generation process. Firstly, control expert generates a video structure. Subsequently, temporal and spatial experts perform the refinement of spatio-temporal details. Spatial and temporal experts work together in refinement stage with our coordinated denoising. By adding consistency initialization, coherence guidance and staggered refinement to ConFiner, ConFiner-Long can generate coherent long videos.
@@
First, we design consistency initialization strategy to promote consistency between segments. The initial noise affects the content of the generated video significantly. To improve the consistency between segments, we first sample a $Noise\_{base}\in\mathbf{R}^{H\times W\times C\times F}$, which is then subjected to frame-wise shuffling to obtain the initial noise for each segment. Sharing base noise enhances the content consistency between segments while shuffling maintains a little randomness.
@@
Although consistency initialization and staggered refinement have ensured content consistency and smooth transitions between segments, if the motion trend of video structures is not coherent, it will be impossible to combine them into a reasonable long video. Thus, we propose a coherent guidance to promote the motion trend of new segment to follow the preceding segment. In video generation, predicted noises affect the direction of generation and determine the motion trend. So we generate each structure one by one, using noises of the previous segments to guide the subsequent structure. Specifically, during the sampling process, we use the gradient of the L2 loss to guide the sampling direction. The L2 loss is calculated between the predicted noise of the current segment and the noise in the previous segment. The guided noise is calculated as follows:
@@
In this way, coherent guidance can make the noise of the two segments similar, which allows the motion trend of the latter segment to inherit that of the previous segment. Additionally, coherence guidance also reduces the pixel distance between noises of two segments, which can help maintain content consistency between segments.
@@
Objective Evaluation